# Position-Sensitive Silicon Photomultiplier Array with Enhanced Position Reconstruction by means of a Deep Neural Network

### v1

@@ Abstract
Single-photon sensitive detectors like Silicon Photomultipliers are widely used in many medical imaging applications. By using detectors with position resolutions, it is possible to build compact photodetector readouts with reduced number of channels, but still preserving position resolution and gamma-rays imaging capabilities. In this work, we present the advantage of using a Deep Neural Networks (DNNs) light position reconstruction applied to a 2x2 array of linearly-graded SiPMs (LG-SiPMs), to minimize the distortions on the reconstructed event maps. Our approach significantly enhances both the resolution and linearity of position detection compared to the nominal reconstruction formula based on the device architecture. Remarkably, the DNN-based reconstruction boosts the number of resolved areas (‘pixels’) by a factor of at least 5.7, allowing a higher level of precision and performance in light detection.
@@ 1 Introduction
Gamma cameras are widely used in medical imaging and other applications, typically based on high-sensitivity photo-detectors that read out scintillator crystals. Examples are Single photon emission computed tomography (SPECT) and positron emission tomography (PET) [AngerCamera, cherry_ch14, Peterson2011, Tsuchimochi_2013, Farnworth2023]. In such applications, position reconstruction is an important part of the acquisition and affect the performance of the overall imaging technique. The scintillating crystal converts gamma-rays into a detectable number of optical photons, which are then converted into an electric signal by the photodetector. The new generation of gamma-ray imaging devices are often employing silicon photomultipliers (SiPMs) as they offer high internal gain, good photo-detection efficiency, good timing resolution, and robustness compared to other detectors like PMTs [Farnworth2023, MedicalSiPM, Gundacker_2020, Acerbi2025, Della_Volpe2025-qd].
@@
Usually, position reconstruction is obtained by having multiple detectors arranged in a 2D resistive network, or using a multi-anode PMT (MA-PMT) with the multi-anode readout with resistive network [Zhang2023, Schaefer1998] such that the position can be inferred from the relative amplitudes of the signals in only 4 channels. This charge-based Center of Gravity (CoG) approach was pioneered in the 50s by Hal O. Anger [AngerCamera]. This approach is advantageous because of the reduction of the number of readout channel, with respect to having to readout one channel per each position. Based on such idea, at FBK (Trento, Italy), a particular type of position-sensitive silicon photomultiplier detector has been developed [Gola2013ANA]. These devices are called linearly-graded SiPMs (LG-SiPMs). Based on such position sensitive technology, it is possible to have large chips (few squared millimeter, up to 1 squared centimeter) with position resolution. These chips can be also arranged in arrays, having a detection module based on 4 elements of linearly-graded SiPMs (LG-SiPMs). Each element joined based on a charge-sharing architecture integrated on the die that has a high resolution, low distortion, and fast output signals [Gola2013ANA, Acerbi2023, Acerbi2024, Gola2020].
@@
In this work, we used a tile of 2x2 elements of LG-SIPM chips, each one with area of about 8x8 mm2, previously characterized in [FBK-LG-SiPM]. The tile covers a nominal chip area of almost 16x16 mm2. The output signals from the LG-SiPMs chips are arranged in a proper configuration, through a ”smart channel” approach, connecting the central ones together, so that the hit position can be inferred with only 6 readout channels [Acerbi2024].
@@
LG-SiPM approach drastically reduces the number of channels to infer the hit position in comparison to an array of SiPMs each having a single readout channel. However, electronic defects and non-uniformities can induce nonlinearity and worsen the position resolution, reducing the number of distinguishable regions and generating distorted images [Acerbi2024].
@@
In this contribution we show a method based on Deep Neural Network (DNN) to minimize these distortion effects thus increasing the granularity of the linearly graded sensor. Neural networks have been investigated on Monte Carlo simulation for event positioning in [Jaliparthi2021DeepRN].
@@ 2 Instrumentation and Methods
We biased the LG-SiPM at 33 V (5 V over-voltage) and amplified the signals of the 6 channels. A blue LED (470 nm peak wavelength) is connected to an optical fiber that is placed in front of the LG-SiPM. Although the light spread should not affect the reconstructed position, since LG-SiPMs are designed to reconstruct the center of gravity of the light beam, we placed the fiber as close as possible to limit the spread of the light. We estimated the light spot size is about 2 mm in diameter.
@@ 2.2 Linear reconstruction
Using the formula for the position reconstruction for an LG-SiPM in [Gola2013ANA] we can derive the relative $x$ and $y$ coordinates for a 2x2 tile of LG-SiPM as:
@@ 2.3 Neural Network reconstruction
The the performance of the DNN for an increased number of layers is shown in section 3.
@@ 2.3.2 Deep Neural Network
To compensate for non-nonlinearities, the model itself should be non-linear. We do this by increasing the number of layers and using non-linear $\tanh$ activation function. The number of units per layer is also increased thus adding more free parameters that captures the sensor non-linear response.
@@
Figure 4: Schematic of the Deep Neural Network architecture used. The DDNs consist of an input layer for the amplitude $Q_{i}$ connected to a series of $N_{\rm{layers}}$ 64-input dense hidden layers with hyperbolic tangent activation functions. The output layer is connected to the last hidden layer by a linear activation function to reconstruct the coordinates.
The inputs are normalized by the total charge for the DNN to be independent of the amount of light (i.e. $Q_{i}\rightarrow\frac{Q_{i}}{\sum_{i=1}^{6}Q_{i}}$). The output layer consists of 2 units dense layer with a linear activation for the $x_{\mathrm{reco}}$ and $y_{\mathrm{reco}}$ reconstructed positions.
@@
The networks were trained on the training sample in batches during 40 epochs using the Adam optimizer. We used the Keras deep learning API written in Python [chollet2015keras].
@@ 3.1 Qualitative results
Reconstructed images of the scan are shown in figure 5 for the splitting technique (A) on the test sample. Each image represents a different number of layers $N_{{\rm layers}}$ used. One can see that the nonlinearities are recovered by the DNN. Quantitative assessment of the DNNs is further discussed in sections 3.3 and 3.4.
@@ 3.3 Resolution and systematic shift
This allows us to evaluate the resolution $\sigma$ (how spread are the reconstructed positions) and average shift $\nu$ of the tested positions (how far is the centroid of the reconstructed positions from the true position). These values were obtained by fitting the distances $d$ to a Rice distribution [rice] as in [raiola2025spatialresolution]. The resolution $\sigma$ represents the noise in the detector that cannot be overcome while the shift $\nu$ is a measure of the displacement that occurs between the true and reconstructed positions that can be learned by the DNNs.
@@ 3.4 Granularity
Another factor of interest is the granularity which relates to the minimum size of distinguishable areas. This quantity estimates the average size of a distinguishable region in the sensor. It is computed as twice the mean of the distance distribution $2\left<d\right$> which corresponds to when two neighboring regions are overlapping and thus are indistinguishable. We observe that the DNN is up to 3.5 times better than the standard reconstruction at distinguishing regions in the sensitive area of the device (see table 1). In other words, for a field of view of 16x16 mm2, the performances observed for the DNN correspond to a maximum of 6400 distinguishable regions while for the standard reconstruction, this corresponds to a maximum of 529 distinguishable regions.
@@
Table 1: Mean resolution $\left<\sigma\right$>, mean shift $\left<\nu\right$>, and granularity of the linear model (top) and DNN (bottom) for the different splitting between the test and train samples. The averaged values were computed among the scanned positions removing the outer rows and columns. The errors correspond to the statistical errors. The three splittings are: random splitting (A), chessboard splitting (B), and random position splitting (C).
@@ 4 Conclusion
We characterized the position reconstruction capabilities of an LG-SiPM comprising only 6 readout channels with a linear reconstruction technique and a DNN. We showed that the device can resolve spots of about 0.67 mm nominally and that non-linearity can be recovered using a trained DNN. The DNN reduces the granularity to about 0.2 mm thus increasing the number of distinguishable regions by a factor 10.24. The intrinsic granularity of LG-SiPMs, further enhanced by the DNN approach described in this work, is significantly finer than the pixelation achievable with current state-of-the-art scintillators used in scintillation cameras. This makes the choice of LG-SiPM suitable for applications requiring sub-millimetric resolution, especially with the implementation of Neural Network-based correction algorithms.

### v2

@@ Abstract
Single-photon sensitive detectors like Silicon Photomultipliers are widely used in many medical imaging applications. By using detectors with position resolutions, it is possible to build compact photodetector readouts with reduced number of channels, but still preserving position resolution and gamma-rays imaging capabilities. In this work, we present the advantage of using a Deep Neural Networks (DNNs) light position reconstruction applied to a 2x2 array of linearly-graded SiPMs (LG-SiPMs), to minimize the distortions on the reconstructed event maps. Our approach significantly enhances both the resolution and linearity of position detection compared to the nominal reconstruction formula based on the device architecture. Remarkably, the DNN-based reconstruction boosts the number of resolved areas (‘pixels’) by a factor of 5.7 to 12.1 (depending the training splitting used) allowing for a higher level of precision and performance in light detection.
@@ 1 Introduction
Gamma cameras are widely used in medical imaging and other applications, typically based on high-sensitivity photo-detectors that read out scintillator crystals. Examples are Single photon emission computed tomography (SPECT) and positron emission tomography (PET) [1, 2, 3, 4, 5]. In such applications, position reconstruction is an important part of the acquisition and affect the performance of the overall imaging technique. The scintillating crystal converts gamma-rays into a detectable number of optical photons, which are then converted into an electric signal by the photodetector. The new generation of gamma-ray imaging devices are often employing silicon photomultipliers (SiPMs) as they offer high internal gain, good photo-detection efficiency, good timing resolution, and robustness compared to other detectors like PMTs [5, 6, 7, 8, 9].
@@
Usually, position reconstruction is obtained by having multiple detectors arranged in a 2D resistive network, or using a multi-anode PMT (MA-PMT) with the multi-anode readout with resistive network [10, 11] such that the position can be inferred from the relative amplitudes of the signals in only 4 channels. This charge-based Center of Gravity (CoG) approach was pioneered in the 50s by Hal O. Anger [1]. This approach is advantageous because of the reduction of the number of readout channel, with respect to having to readout one channel per each position. Based on such idea, at FBK (Trento, Italy), a particular type of position-sensitive silicon photomultiplier detector has been developed [12]. These devices are called linearly-graded SiPMs (LG-SiPMs). Based on such position sensitive technology, it is possible to have large chips (few squared millimeter, up to 1 squared centimeter) with position resolution. These chips can be also arranged in arrays, having a detection module based on 4 elements of linearly-graded SiPMs (LG-SiPMs). Each element joined based on a charge-sharing architecture integrated on the die (see figure 1-left) that has a high resolution, low distortion, and fast output signals [12, 13, 14, 15].
Figure 1: Left: Scheme showing a portion of a single LG-SiPM chip’s micro-cells connected to the resistive/capacitive network allowing to reconstruct the hit position from 4 channels. Taken from [14]. Right: Connection scheme of the 2x2 LG-SiPM elements arranged in a ”smart channel” approach. The position can be reconstructed with equation 1. Adapted from [14].
@@
In this work, we used a tile of 2x2 elements of LG-SiPM chips, each one with area of about 8x8 mm2, previously characterized in [16]. The tile covers a nominal chip area of almost 16x16 mm2. The output signals from the LG-SiPMs chips are arranged in a proper configuration, through a ”smart channel” approach, connecting the central ones together, so that the hit position can be inferred with only 6 readout channels as shown in figure 1-right [14].
@@
LG-SiPM approach drastically reduces the number of channels to infer the hit position in comparison to an array of SiPMs each having a single readout channel. However, electronic defects and non-uniformities can induce non-linearity and worsen the position resolution, reducing the number of distinguishable regions and generating distorted images [14].
@@
In this contribution we show a method based on Deep Neural Network (DNN) to minimize these distortion effects thus increasing the granularity of the linearly graded sensor. Neural networks have been investigated on Monte Carlo simulation for event positioning in [17].
@@ 2 Instrumentation and Methods
We biased the LG-SiPM at 33 V (5 V over-voltage) and amplified the signals of the 6 channels. A blue LED (470 nm peak wavelength) is connected to an optical fiber that is placed in front of the LG-SiPM. Although the light spread should not affect the reconstructed position, since LG-SiPMs are designed to reconstruct the center of gravity of the light beam, we placed the fiber as close as possible to limit the spread of the light. We estimated that the light spot size is about 2 mm in diameter.
@@ 2.2 Linear reconstruction
Using the formula for the position reconstruction for an LG-SiPM in [12] we can derive the relative $x$ and $y$ coordinates for a 2x2 tile of LG-SiPM as:
@@
The channel number $Q_{i}$ can be identified from figure 1-right.
@@ 2.3 Neural Network reconstruction
The performance of the DNN for an increased number of layers is shown in section 3.
@@ 2.3.2 Deep Neural Network
To compensate for non-linearities, the model itself should be non-linear. We do this by increasing the number of layers and using non-linear hyperbolic tangent activation functions. The number of units per layer is also increased thus adding more free parameters that captures the sensor non-linear response.
@@
Figure 5: Schematic of the Deep Neural Network architecture used. The DNNs consist of an input layer for the amplitude $Q_{i}$ connected to a series of $N_{\rm{layers}}$ 64-input dense hidden layers with hyperbolic tangent activation functions. The output layer is connected to the last hidden layer by a linear activation function to reconstruct the coordinates.
The dataset was divided into independent training and testing samples as explained in 2.1. The training sample was used to optimize the network weights and biases, while the validation sample was used to evaluate the performance of the networks (see section 3).
Before training, the input $Q_{i}$ were normalized by the total collected charge in order to remove the dependence on the absolute light yield:
$Q_{i}\rightarrow\frac{Q_{i}}{\sum_{i=1}^{6}Q_{i}}.$ (7)
This normalization stabilizes the training and ensures that the network learns from the relative charge distribution rather than from the absolute signal amplitudes.
The output layer consists of 2 units dense layer with a linear activation for the $x_{\mathrm{reco}}$ and $y_{\mathrm{reco}}$ reconstructed positions.
@@
The networks were trained in batches during 40 epochs using the Adam optimizer [18] with an initial learning rate $\alpha=10^{-3}$, exponential decay rates $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=10^{-7}$.
The model weights were initialized using the Glorot uniform initialization, which is well suited for hyperbolic tangent activation functions and helps maintain stable gradients during training [19]. The model biases were initialized to zero.
Training convergence was verified by monitoring both the training and validation losses as a function of epoch number (see section 3.2).
All models were implemented using the Keras deep learning API written in Python [20].
@@ 3.1 Qualitative results
Reconstructed images of the scan are shown in figure 6 for the splitting technique (A) on the test sample. Each image represents a different number of layers $N_{{\rm layers}}$ used. One can see that the non-linearities are recovered by the DNN. The reconstructed images from the linear reconstruction presented in section 2.2 can be found in [21].
Large position reconstruction shift can be observed at the gaps of the SiPM tiles, borders of the LG-SiPMs and in some rows or columns of particular SiPM tiles systematically. Although these discrepancies are not completely overcome by the DNN we observe that the reconstruction error is improved. These systematic shift in the position errors could be explained by the imperfections in the linearly-graded resistive network or the quality cut which inherently cuts events at the borders and in the gaps thus reducing sample to train events in these regions.
Quantitative assessment of the DNNs is further discussed in sections 3.3 and 3.4.
@@ 3.3 Resolution and systematic shift
This allows us to evaluate the resolution $\sigma$ (how spread are the reconstructed positions) and average shift $\nu$ of the tested positions (how far is the centroid of the reconstructed positions from the true position). These values were obtained by fitting the distances $d$ to a Rice distribution [22] as in [21]. The resolution $\sigma$ represents the noise in the detector that cannot be overcome while the shift $\nu$ is a measure of the displacement that occurs between the true and reconstructed positions that can be learned by the DNNs.
@@ 3.4 Granularity
Another factor of interest is the granularity which relates to the minimum size of distinguishable areas. This quantity estimates the average size of a distinguishable region in the sensor. It is computed as twice the mean of the distance distribution $2\left<d\right$> which corresponds to when two neighboring regions are overlapping and thus are indistinguishable. Results of the granularity are reported in table 1.
3.5 Number of resolvable regions
Additionally to the granularity, one can compute the effective number of distinguishable regions for each method investigated. This number is given by the ratio of the device sensitive area $A=16\times 16~{\rm mm^{2}}$ to the granularity squared:
$N=\frac{A}{4\left<d\right>^{2}}.$ (9)
We observe that the number of distinguishable region is up to $12.1$ times greater from the DNN model compared to the linear model when considering the splitting (A). For the other splitting methods the improvement factors are $9$ (splitting (B)) and $5.7$ (splitting (C)) In other words, for a field of view of 16x16 mm2, the performances observed for the DNN correspond to a maximum of $6530$ distinguishable regions while for the standard reconstruction, this corresponds to a maximum of $541.5$ distinguishable regions.
@@ 3.5 Number of resolvable regions
N. of resolvable regions NN $6530.0\pm 6.6$ $4894.0\pm 4.3$ $3182.9\pm 2.2$
Table 1: Mean resolution $\left<\sigma\right$>, mean shift $\left<\nu\right$>, granularity and number of resolvable regions of the linear model (top) and DNN (bottom) for the different splitting between the test and train samples. The averaged values were computed among the scanned positions removing the outer rows and columns. The errors correspond to the statistical errors. The three splittings are: random splitting (A), chessboard splitting (B), and random position splitting (C).
@@ 4 Conclusion
We characterized the position reconstruction capabilities of an LG-SiPM comprising only 6 readout channels with a linear reconstruction technique and a DNN. We showed that the device has a granularity of about 0.69 mm nominally and that non-linearity can be recovered using a trained DNN granularity to about 0.20 mm when trained on a uniform random splitting of the data. In this case, the DNN reduces the number of distinguishable regions by a factor 12.1. The intrinsic granularity of LG-SiPMs, further enhanced by the DNN approach described in this work, is significantly finer than the pixelation achievable with current state-of-the-art scintillators used in scintillation cameras. This makes the choice of LG-SiPM suitable for applications requiring sub-millimetric resolution, especially with the implementation of Neural Network-based correction algorithms.
The present results are based on LED‑scan, we could study the improvements with real scintillation light used for radiation detection. The implications of scintillator size and optical spread, as well as the effect of microcell occupancy and potential saturation may modify the absolute position resolution when using real scintillators. However this do not affect the validity of the comparative analysis between the nominal (linear) reconstruction and the DNN approaches when considering only the intrinsic LG-SiPM capabilities. Future work will include measurements with scintillator‑coupled configurations to experimentally confirm the robustness of the DNN for radiation detection applications such as gamma-ray imaging. However it might be challenging, in a scintillator‑coupled system, to control the exact position and direction of incoming gamma‑rays due to the stochastic nature of radioactive decay, which limits the precision of beam‑based validation. As a consequence, precise ground‑truth position labels cannot be obtained experimentally, making LED‑based scans and Monte Carlo simulations he only practical way to perform supervised studies of reconstruction accuracy. This motivates the light‑only calibration strategy adopted in this work.